How many pounds of raisins worth $3 per pound must be fixed with 10 pounds of peanuts worth $2.40 per pound to produce a mixture worth $2.80 per pound
Answer by Gersid(33)

Try this:
Let x = the required number of pounds of raisins @ $3.00 per pound.
$3.00(x)+$2.40(10) = $2.80(10+x) Simplify and solve for x.
3x+24 = 28+2.8x Subtract 2.8x from both sides.
0.2x+24 = 28 Subtract 24 from both sides.
0.2x = 4 Finally, divide both sides by 0.2
x = 20
You will need 20 pounds of raisins.